# Theoretical Bounds for Stable In-Context Learning

### v1

@@ 7 Ethics Statement
We used large language models (LLMs) in two limited ways. First, for aid or polish writing, an LLM service (provider redacted for anonymous review) helped with grammar, style, and tightening of sentences. All technical content, problem formulations, theorems, proofs, algorithmic designs, model implementations, experiments, and final claims were authored and verified by the authors. Second, for retrieval and discovery, we used LLM-assisted search to expand keyword queries and to draft short summaries of candidate papers from public abstracts. Inclusion decisions and all citations were made only after the authors read the original sources; no autogenerated references were accepted. No private or proprietary data were provided to the LLM, and prompts contained only our own drafts or publicly available text. This use does not affect reproducibility: the released code, data, and results do not depend on any LLM outputs.
Appendix B Robustness Extensions
@@
In the main text, our theoretical guarantees are derived under the nominal assumption (A1), namely that the representation $\phi(x)$ is fixed and sub-Gaussian with covariance $\Sigma$. Here we discuss how the results extend when assumptions (A2)–(A4) are considered.
B.1 Robustness to Representation Drift (A2)
We first allow the representation to vary with context index $t$, modeled as
@@ B.1 Robustness to Representation Drift (A2)
where $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$ uniformly in $x,t$. Let $\Sigma_{t}=\mathbb{E}[\phi_{t}(x)\phi_{t}(x)^{\top}]$ and denote the averaged covariance by $\bar{\Sigma}=\tfrac{1}{T}\sum_{t=1}^{T}\Sigma_{t}$.
@@
Robustness under bounded drift
Let $K_{\mathrm{nominal}}$ be the sample-size lower bound under (A1), depending on $\Sigma$. Under (A2), the empirical covariance $\hat{\Sigma}_{K}$ concentrates around $\bar{\Sigma}$, and for any $\delta\in(0,\lambda_{\min}(\bar{\Sigma}))$, with probability at least $1-\xi$ we have
$\lambda_{\min}(\hat{\Sigma}_{K})\;\geq\;\delta-c_{1}\varepsilon,$
provided
$K\;\;\geq\;\;K_{\mathrm{nominal}}\;+\;c_{2}\cdot\frac{\varepsilon^{2}\|\Sigma\|}{(\lambda_{\min}(\Sigma)-\delta)^{2}}\,$,
where $c_{1},c_{2}$ are absolute constants.
Proof sketch.
Write $\phi_{t}(x)=\phi(x)+\Delta_{t}(x)$ and expand $\Sigma_{t}=\Sigma+\mathbb{E}[\phi(x)\Delta_{t}(x)^{\top}]+\mathbb{E}[\Delta_{t}(x)\phi(x)^{\top}]+\mathbb{E}[\Delta_{t}(x)\Delta_{t}(x)^{\top}]$. The last three terms are bounded in operator norm by $O(\varepsilon\|\Sigma\|^{1/2}+\varepsilon^{2})$. Applying Weyl’s inequality gives $\lambda_{\min}(\Sigma_{t})\geq\lambda_{\min}(\Sigma)-O(\varepsilon)$. Matrix Bernstein then yields concentration of $\hat{\Sigma}_{K}$ around $\bar{\Sigma}$ with the same rate as before, plus an additive slack proportional to $\varepsilon$. The additional sample-size term arises from controlling the variance proxy $v$ with $\varepsilon$-dependent terms. ∎
B.2 Relaxing Tail Assumptions (A3)
@@
B.3 Weak Dependence among Features (A4)
When $\{\phi_{t}(x)\}$ are not independent across $t$, we apply matrix concentration for martingale-difference or $\alpha$-mixing sequences (e.g., Tropp 2011, Paulin 2015). In this case, the effective sample size is reduced: if the mixing coefficient is $\rho\in[0,1)$, then the bound takes the same form as in Theorem B.1, with $K$ replaced by $K/(1+\rho)$. Thus moderate dependence inflates the sample complexity by at most a constant factor.
B.4 Summary
@@
Appendix C Extensions to Dependent Features
@@
$\Sigma,\ \|\Sigma\|$ Covariance and operator norm (Sec. 3)
$r_{\mathrm{eff}}=\mathrm{tr}(\Sigma)/\|\Sigma\|$ Effective rank (Sec. 3; Appx. G.2)
$\sigma$ Sub-Gaussian parameter (Appx. 3.2; Appx. G.3)
@@
$C$ Constant in operator-norm deviation equation(9)
$C^{\prime},C^{\prime\prime}$ Constants in plug-in lower bounds equation(10), (13)
$B_{K}$ Truncation threshold (equation(6))
$\lambda_{q}$ $q$-quantile eigenvalue surrogate (Appx. F)
@@
C.1 Martingale-difference setting
Suppose $\{\phi_{t}(x)\}$ forms a martingale-difference sequence with respect to a natural filtration $\mathcal{F}_{t}$ (e.g., conditioned on past tokens). That is, $\mathbb{E}[\phi_{t}(x)\mid\mathcal{F}_{t-1}]=0$. Then the sequence of covariance deviations

### v2

@@ A.7 Connection to ICL (in-context learning) and experimental suggestions
We used large language models (LLMs) in two limited ways. First, for aid or polish writing, an LLM service (provider redacted for anonymous review) helped with grammar, style, and tightening of sentences. All technical content, problem formulations, theorems, proofs, algorithmic designs, model implementations, experiments, and final claims were authored and verified by the authors. Second, for retrieval and discovery, we used LLM-assisted search to expand keyword queries and to draft short summaries of candidate papers from public abstracts. Inclusion decisions and all citations were made only after the authors read the original sources; no autogenerated references were accepted. No private or proprietary data were provided to the LLM, and prompts contained only our own drafts or publicly available text.
Appendix C Experimental Reproducibility Specification
This appendix provides an executable specification of our experimental pipeline.
C.1 Datasets and splits
We evaluate (i) text classification tasks: SST-2, AGNews, Yahoo! Answers (Topics), and TREC-6; (ii) multi-step reasoning: GSM8K (and SVAMP in extended ablations); and (iii) retrieval-augmented generation (RAG): HotpotQA (and NaturalQuestions in extended ablations). Unless otherwise stated, for each task we use the standard public train/test split. Demonstrations are sampled from the training split; evaluation queries are drawn from the test split. For calibration (when used), we draw a small validation subset from the training split only (never from the test split) to avoid test leakage.
C.2 Prompt template and label mapping
All ICL prompts follow the same canonical format, with $K$ demonstrations followed by a query:
Task: {task_name}
You will be given an input and you must output the label. Output only the label token.
Input: {x_1}
Label: {y_1}
Input: {x_K}
Label: {y_K}
Input: {x_q}
Label:
Single-token labels.
To reduce output ambiguity, we use single-token labels whenever possible. For binary tasks (e.g., SST-2), we map labels to yes / no. For multi-class tasks, we map each class to a distinct single token (e.g., a digit token 0, 1, …), and we provide an explicit class-to-token mapping in the prompt header. When a dataset’s native label string is multi-token, we report results under both the native label and the single-token remapping in the corresponding ablation.
C.3 Demonstration selection and retrieval
Unless an ablation specifies otherwise, we use cosine-similarity nearest-neighbor retrieval (Cosine-NN) to form demonstrations for each query:
  • Embed each candidate training input $x$ and query $x_{q}$ into a vector space (Section C.4).
  • Rank training candidates by cosine similarity to $x_{q}$ and select the top-$K$ as demonstrations.
In the retrieval ablation, we additionally compare: (i) random sampling; and (ii) class-balanced sampling (with equal per-class quota when feasible). Throughout, “retrieval” in this subsection refers to selecting demonstrations; evidence retrieval for RAG is specified separately in Appendix K.
C.4 Encoders and representations
Encoders.
We use sentence-level encoders listed in Table 1 (e.g., all-mpnet-base-v2, all-MiniLM-L6-v2, paraphrase-MiniLM-L12-v2). For each input text $x$, we compute a fixed embedding $e(x)\in\mathbb{R}^{d}$ using the encoder’s default pooling (e.g., mean pooling over token embeddings followed by normalization if provided by the encoder).
Second-moment construction.
Given $K$ demonstrations selected for a query, we form the empirical second-moment matrix $\widehat{\Sigma}_{K}:=\tfrac{1}{K}\sum_{i=1}^{K}e(x_{i})e(x_{i})^{\top}$ from their encoder embeddings (using the same normalization/preprocessing across all experiments). When a ridge-shifted proxy is used, we report results with $\widehat{\Sigma}_{K,\rho}:=\widehat{\Sigma}_{K}+\rho I$ where $\rho$ is computed once from the pilot matrix as $\rho=\min\{\epsilon_{\rho}\cdot\mathrm{tr}(\widehat{\Sigma}_{K_{0}})/d,\ \delta/2\}$ and is then held fixed for all $K$ within that run. The coefficient $\epsilon_{\rho}$ is fixed across runs; calibration (when enabled) tunes only $(q,\alpha)$ unless explicitly stated.
Generator internal representations.
Most commercial LLM APIs do not expose internal hidden states. Accordingly, our diagnostics focus on observable quantities (accuracy curves, knee estimates, and spectral proxies computed from external encoders).
C.5 Commercial API generators and decoding hyperparameters
Between 12.18.2025 and 1.15.2026, we evaluate six large-scale commercial API models latest version: OpenAI gpt-5.2-chat-latest; Anthropic claude-sonnet-4-5; Gemini gemini-3-flash-preview; DeepSeek deepseek-reasoner; Kimi kimi-k2-thinking; Qwen qwen-max-latest.
Deterministic decoding.
To isolate instability due to demonstrations rather than sampling, we use deterministic decoding whenever supported: temperature =0=0, top-$p=1.0$ (and equivalent provider parameters), and greedy/argmax decoding. We set max output to $1$ label token under the single-token label mapping; in the multi-token ablation we set max output to the minimal required tokens and stop on newline/EOS.
Request logging.
For each request we log: provider, model ID/alias, API endpoint/base_url, request timestamp (UTC), decoding parameters, and the raw completion. Because provider-side models can evolve behind stable aliases, we report exact run windows and store the returned model identifier when available.
C.6 Random seeds and determinism
All reported results average over $5$ random seeds. We use seeds $\{0,1,2,3,4\}$ and set them for Python random, NumPy, and PyTorch. For GPU runs we enable deterministic kernels when available and log the exact library versions in the run metadata. The only stochasticity remaining under our default protocol comes from (i) data subsampling (when used) and (ii) bootstrapping for confidence intervals.
Reporting variance across seeds.
For each configuration, we compute the ratio $r_{s}:=K^{*}_{s}/\widehat{K}_{\mathrm{knee},s}$ separately for each seed $s\in\{0,1,2,3,4\}$ (where $\widehat{K}_{\mathrm{knee},s}$ is the seed-specific knee estimated on the test split), and report the mean $\frac{1}{5}\sum_{s}r_{s}$ in the main tables. We additionally report the standard deviation $\sqrt{\frac{1}{4}\sum_{s}(r_{s}-\bar{r})^{2}}$ in the released artifact (along with per-seed knees and accuracies), since it is often more informative than reporting variance only on accuracies.
C.7 K-grid, knee-point estimator, and confidence intervals
$K$ grid.
We evaluate accuracy on a discrete grid $\mathcal{K}$ of prompt lengths: $\mathcal{K}=\{0,1,2,4,8,16,32,64,128\}$, truncated as needed to respect each model’s context window.
Knee-point estimator.
Given the curve $\{(K,\mathrm{ACC}(K))\}_{K\in\mathcal{K}}$, we fit a two-segment piecewise linear model. For each candidate split index $t$ (with at least two points on each side), we fit two least-squares lines on $\{K\leq t\}$ and $\{K>t\}$ and compute the sum of squared errors (SSE). We define $\widehat{K}_{\mathrm{knee}}$ as the $K$ at the split minimizing SSE (ties broken by choosing the smaller $K$ for conservatism).
Bootstrap confidence intervals.
We compute a $95\%$ percentile bootstrap interval for $\widehat{K}_{\mathrm{knee}}$ using $B=200$ resamples. Each resample draws evaluation queries with replacement from the test set, recomputes $\mathrm{ACC}(K)$ for all $K\in\mathcal{K}$ (with demonstrations formed under the same fixed protocol), and recomputes $\widehat{K}_{\mathrm{knee}}$. We report the 2.5/97.5 percentiles over the $B$ bootstrapped knees.
Sanity checks.
To verify that the two-segment fit is not spuriously triggered by smooth saturation or noise, we inspect representative fitted curves and confirm that the selected breakpoint aligns with a visible transition region; we include example curves and fitted segments in the supplementary material.
C.8 Calibration protocol (no test leakage)
When calibration is enabled (choosing $q$ for $\lambda_{q}$ and fitting the global scale factor $\alpha$), we tune hyperparameters on a small calibration subset drawn from training data only and keep the test split untouched.
Task-level split.
For each encoder–generator pair, we select a small set of tasks $\mathcal{T}_{\mathrm{cal}}$ for calibration and a disjoint set $\mathcal{T}_{\mathrm{eval}}$ for final reporting. The selection is fixed before running any test evaluation.
Within-task data split.
For each task, we use the training split to form demonstrations and (when needed) to build the calibration curves. All reported accuracies, knee-points, and final ratios are computed on the task’s test split.
Fitting $(q,\alpha)$ from measured knee-points.
For each $t\in\mathcal{T}_{\mathrm{cal}}$, we estimate $\widehat{K}_{\mathrm{knee}}^{(t)}$ using the piecewise-linear method described above, and compute $\widehat{K}^{\mathrm{raw}}_{t}$ from (27) using the pilot matrix $\widehat{\Sigma}_{K_{0}}$. We choose $q\in[0.05,0.2]$ by minimizing the median absolute log-error on $\mathcal{T}_{\mathrm{cal}}$:
$q^{\star}\in\arg\min_{q\in[0.05,0.2]}\operatorname{median}_{t\in\mathcal{T}_{\mathrm{cal}}}\left|\log\widehat{K}^{\mathrm{raw}}_{t}(q)-\log\widehat{K}_{\mathrm{knee}}^{(t)}\right|.$
Then we set $\alpha$ by the median ratio fit (29) with $q=q^{\star}$. Finally, we freeze $(q^{\star},\alpha)$ and evaluate on $\mathcal{T}_{\mathrm{eval}}$ without further tuning.
Appendix D Robustness Extensions
@@
In the main text, our theoretical guarantees are derived under the nominal assumption (A1), namely that the representation $\phi(x)$ is fixed and sub-Gaussian with second-moment $\Sigma$. Here we discuss how the results extend when assumptions (A2)–(A4) are considered.
D.1 Robustness to Representation Drift (A2)
We consider an independent but non-identically distributed (i.n.i.d.) sequence across the $K$ in-context demonstrations. For each context index $t=1,\dots,K$, let $x_{t}\sim\mathcal{D}$ be drawn independently and define
@@ D.1 Robustness to Representation Drift (A2)
where the drift satisfies the uniform bound $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$ for all $x,t$. Let
$X_{t}:=\phi_{t}(x_{t})\phi_{t}(x_{t})^{\top},\qquad\Sigma_{t}:=\mathbb{E}[X_{t}],\qquad\bar{\Sigma}:=\frac{1}{K}\sum_{t=1}^{K}\Sigma_{t},\qquad\widehat{\Sigma}_{K}:=\frac{1}{K}\sum_{t=1}^{K}X_{t}.$
The goal is to control $\lambda_{\min}(\widehat{\Sigma}_{K})$ via concentration of $\widehat{\Sigma}_{K}$ around $\bar{\Sigma}$, and then relate $\bar{\Sigma}$ back to the nominal $\Sigma:=\mathbb{E}[\phi(x)\phi(x)^{\top}]$.
Lemma D.1 (Second-moment shift under bounded drift).
Under $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$, for every $t$,
$\|\Sigma_{t}-\Sigma\|_{\mathrm{op}}\;\leq\;2\varepsilon\,\|\Sigma\|_{\mathrm{op}}^{1/2}+\varepsilon^{2},\qquad\text{and hence}\qquad\|\bar{\Sigma}-\Sigma\|_{\mathrm{op}}\;\leq\;2\varepsilon\,\|\Sigma\|_{\mathrm{op}}^{1/2}+\varepsilon^{2}.$
Proof.
Expanding $\Sigma_{t}=\mathbb{E}[(\phi+\Delta_{t})(\phi+\Delta_{t})^{\top}]$ yields
$\Sigma_{t}-\Sigma=\mathbb{E}[\phi\Delta_{t}^{\top}]+\mathbb{E}[\Delta_{t}\phi^{\top}]+\mathbb{E}[\Delta_{t}\Delta_{t}^{\top}].$
Since $\Delta_{t}\Delta_{t}^{\top}\preceq\|\Delta_{t}\|_{2}^{2}I\preceq\varepsilon^{2}I$, we have $\|\mathbb{E}[\Delta_{t}\Delta_{t}^{\top}]\|_{\mathrm{op}}\leq\varepsilon^{2}$. For the cross term, use the operator Cauchy–Schwarz inequality:
$\|\mathbb{E}[\phi\Delta_{t}^{\top}]\|_{\mathrm{op}}\;\leq\;\big\|\mathbb{E}[\phi\phi^{\top}]\big\|_{\mathrm{op}}^{1/2}\,\big\|\mathbb{E}[\Delta_{t}\Delta_{t}^{\top}]\big\|_{\mathrm{op}}^{1/2}\;\leq\;\|\Sigma\|_{\mathrm{op}}^{1/2}\cdot\varepsilon.$
The same bound holds for $\|\mathbb{E}[\Delta_{t}\phi^{\top}]\|_{\mathrm{op}}$, proving the claim. ∎
Variance-proxy derivation (i.n.i.d.).
For matrix Bernstein in the i.n.i.d. setting, we center each summand by its own mean:
$Y_{t}:=X_{t}-\Sigma_{t},\qquad\mathbb{E}[Y_{t}]=0,$
and the variance parameter is
$v:=\left\|\sum_{t=1}^{K}\mathbb{E}[Y_{t}^{2}]\right\|_{\mathrm{op}}\;\leq\;\sum_{t=1}^{K}\left\|\mathbb{E}[Y_{t}^{2}]\right\|_{\mathrm{op}}\;=\;K\cdot\max_{t}\left\|\mathbb{E}[Y_{t}^{2}]\right\|_{\mathrm{op}}.$
We now bound $\|\mathbb{E}[Y_{t}^{2}]\|_{\mathrm{op}}$ explicitly under drift.
Lemma D.2 (Variance proxy under bounded drift).
Let $Y_{t}=X_{t}-\Sigma_{t}$ with $X_{t}=\phi_{t}\phi_{t}^{\top}$ and $\|\Delta_{t}\|_{2}\leq\varepsilon$. Then for every $t$,
$\left\|\mathbb{E}[Y_{t}^{2}]\right\|_{\mathrm{op}}$ $\leq 2\,\mathbb{E}\|X_{t}\|_{\mathrm{op}}^{2}+2\,\|\Sigma_{t}\|_{\mathrm{op}}^{2}$
$=2\,\mathbb{E}\|\phi_{t}\|_{2}^{4}+2\,\|\Sigma_{t}\|_{\mathrm{op}}^{2}$
$\leq 16C_{1}\sigma^{4}\;+\;48\,\varepsilon^{2}\,\mathrm{tr}(\Sigma)\;+\;16\varepsilon^{4}\;+\;2\big(\|\Sigma\|_{\mathrm{op}}+2\varepsilon\|\Sigma\|_{\mathrm{op}}^{1/2}+\varepsilon^{2}\big)^{2},$
where $C_{1}$ is the fourth-moment constant from (A1) (i.e., $\mathbb{E}\|\phi\|_{2}^{4}\leq C_{1}\sigma^{4}$).
Proof.
Using the noncommutative square inequality $(A-B)^{2}\preceq 2A^{2}+2B^{2}$ for symmetric $A,B$ (cf. (17) in the main text), we have
$Y_{t}^{2}=(X_{t}-\Sigma_{t})^{2}\preceq 2X_{t}^{2}+2\Sigma_{t}^{2}.$
Taking expectation and operator norm, and using $\|\mathbb{E}[M]\|_{\mathrm{op}}\leq\mathbb{E}\|M\|_{\mathrm{op}}$, gives
$\|\mathbb{E}[Y_{t}^{2}]\|_{\mathrm{op}}\leq 2\,\mathbb{E}\|X_{t}^{2}\|_{\mathrm{op}}+2\|\Sigma_{t}^{2}\|_{\mathrm{op}}=2\,\mathbb{E}\|X_{t}\|_{\mathrm{op}}^{2}+2\|\Sigma_{t}\|_{\mathrm{op}}^{2},$
since $X_{t}\succeq 0$ implies $\|X_{t}^{2}\|_{\mathrm{op}}=\|X_{t}\|_{\mathrm{op}}^{2}$. Next, $\|X_{t}\|_{\mathrm{op}}=\|\phi_{t}\|_{2}^{2}$, hence $\mathbb{E}\|X_{t}\|_{\mathrm{op}}^{2}=\mathbb{E}\|\phi_{t}\|_{2}^{4}$.
To bound $\mathbb{E}\|\phi_{t}\|_{2}^{4}$, use $\|\phi_{t}\|_{2}\leq\|\phi\|_{2}+\|\Delta_{t}\|_{2}\leq\|\phi\|_{2}+\varepsilon$ and the inequality $(a+b)^{4}\leq 8a^{4}+24a^{2}b^{2}+8b^{4}$ for $a,b\geq 0$:
$\|\phi_{t}\|_{2}^{4}\leq 8\|\phi\|_{2}^{4}+24\varepsilon^{2}\|\phi\|_{2}^{2}+8\varepsilon^{4}.$
Taking expectation and using $\mathbb{E}\|\phi\|_{2}^{4}\leq C_{1}\sigma^{4}$ and $\mathbb{E}\|\phi\|_{2}^{2}=\mathrm{tr}(\Sigma)$ yields
$\mathbb{E}\|\phi_{t}\|_{2}^{4}\leq 8C_{1}\sigma^{4}+24\varepsilon^{2}\mathrm{tr}(\Sigma)+8\varepsilon^{4}.$
Finally, by Lemma D.1, $\|\Sigma_{t}\|_{\mathrm{op}}\leq\|\Sigma\|_{\mathrm{op}}+2\varepsilon\|\Sigma\|_{\mathrm{op}}^{1/2}+\varepsilon^{2}$. Combining the bounds proves the claim. ∎
@@
Theorem D.3 (Robustness under bounded drift (with explicit variance proxy)).
Under (A2), apply matrix Bernstein to the independent centered matrices $Y_{t}=X_{t}-\Sigma_{t}$. Let $v$ be the Bernstein variance parameter and let $B$ be any almost-sure bound on $\|Y_{t}\|_{\mathrm{op}}$ after truncation (as in the main proof technique). Then for any $\eta>0$,
$\Pr\!\left(\left\|\widehat{\Sigma}_{K}-\bar{\Sigma}\right\|_{\mathrm{op}}\geq\eta\right)\;\leq\;d\cdot\exp\!\left(-\frac{K\eta^{2}/2}{v+B\eta/3}\right).$
Moreover, Lemma D.2 provides an explicit $\varepsilon$-dependent upper bound on $v$ (via $v\leq K\max_{t}\|\mathbb{E}[Y_{t}^{2}]\|_{\mathrm{op}}$), and Lemma D.1 relates $\bar{\Sigma}$ back to $\Sigma$. Consequently, whenever $\eta<\lambda_{\min}(\bar{\Sigma})-\delta$, the event $\|\widehat{\Sigma}_{K}-\bar{\Sigma}\|_{\mathrm{op}}\leq\eta$ implies
$\lambda_{\min}(\widehat{\Sigma}_{K})\geq\lambda_{\min}(\bar{\Sigma})-\eta\geq\delta.$
D.2 Relaxing Tail Assumptions (A3)
@@
D.3 Weak Dependence among Features (A4)
When $\{\phi_{t}(x)\}$ are not independent across $t$, we apply matrix concentration for martingale-difference or $\alpha$-mixing sequences . In this case, the effective sample size is reduced: if the mixing coefficient is $\rho\in[0,1)$, then the bound takes the same form as in Theorem D.3, with $K$ replaced by $K/(1+\rho)$. Thus moderate dependence inflates the sample complexity by at most a constant factor.
D.4 Summary
@@
Appendix E Extensions to Dependent Features
@@
$\Sigma,\ \|\Sigma\|$ Covariance and operator norm
$r_{\mathrm{eff}}=\mathrm{tr}(\Sigma)/\|\Sigma\|$ Effective rank
$\sigma$ Sub-Gaussian parameter
@@
$C$ Constant in operator-norm deviation
$C^{\prime},C^{\prime\prime}$ Constants in plug-in lower bounds
$B_{K}$ Truncation threshold
$\lambda_{q}$ $q$-quantile eigenvalue surrogate
@@
E.1 Martingale-difference setting
Suppose $\{\phi_{t}(x)\}$ forms a martingale-difference sequence with respect to a natural filtration $\mathcal{F}_{t}$ (e.g., conditioned on past tokens). That is, $\mathbb{E}[\phi_{t}(x)\mid\mathcal{F}_{t-1}]=0$. Then the sequence of second-moment deviations